OpenStudy (anonymous):

Use the distributive property to simplify 8(3x + 1). A. 24x B. 24x + 1 C. 24x + 8 D. 3x + 8

OpenStudy (anonymous):

it's c

OpenStudy (anonymous):

can you tell me how you got the answer

OpenStudy (anonymous):

multiply 3x by 8 then +1 by 8

OpenStudy (anonymous):

8(3x+1) =8*3x+8*1 =24x+8

OpenStudy (anonymous):

thank you both

OpenStudy (anonymous):

yw

OpenStudy (anonymous):

c